Ascorbic acid is a diprotic acid (Ka1 6.9x10-5) and (ka2 = 2.7x10-12). What is the pH of

a solution that contains 5.0 mg of acid per milliter solution.


1
Expert's answer
2022-01-06T22:03:01-0500

Molar mass of ascorbic acid "=176.12g"

Concentration"=\frac {5×10^{-3}}{176.12}×10^3"


"=0.02839M"



"ka_1\ggg\>ka_2"


"\implies" First dissociation produces virtually all of "H_3O^+"


"H_2C_6H_6O_{6(aq)}+H_2O_{(l)}\rightleftharpoons HC_6H_6O_6^-(aq)+H_3O^+(aq)"





"\begin{matrix}
 ICE&H_2C_6H_6O_6&HC_6H_6O_6^-&H_3O^+\\
 I&0.02839&0&0 \\
C&-x&x&x\\
E&0.02839-x&x&x
\end{matrix}"




"6.9×10^{-5}=\frac{(x)(x)}{(0.02839-x)}\>\>\>x\lll0.02839"


"\therefore\>x^2=1.959×10^{-6}"

"x=1.4\u00d710^{-3}"



"PH=-" log "(1.4\u00d710^{-3})"


"=2.854"